Two electrons approaches each other with equal speeds 0.9c. Find the speed of one electron w.r.t. the other.


1
Expert's answer
2021-04-08T09:00:48-0400

Explanations & Calculations


  • Since the speeds are relativistic, only the relativistic velocity addition can be used.
  • Using the equation "\small V_{relative}=\large \frac{U+v}{1+\large\frac{U\times v}{C^2}}" where "\small U" is the velocity of a frame S, "v" the velocity of any arbitrary particle relative to the frame S & "\small V" is the relative velocity of the particle relative to a stationary frame P.
  • For the ease stop one frame & put its velocity on the other
  • Then the knowns are "\small U=\rightarrow 0.9C" and "\small v=\rightarrow 0.9C"
  • Then,

"\qquad\qquad
\begin{aligned}
\small V_{relative}&=\small \frac{0.9C+0.9C}{1+\large\frac{0.9C\times0.9C}{C^2}}\\
&=\small \frac{1.8C}{1.81}\\
& =\small\bold{0.9945C}
&\small 
\end{aligned}"

  • Each electron sees the other electron approaching with a velocity of 0.9945C